THE v_{t} = 3 STATE OF METHANOL

\underline{JOHN C. PEARSON}, BRIAN J. DROUIN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109; ROGIER BRAAKMAN, GEOFFREY A. BLAKE, Planetary Sciences, California Institute of Technology, Mail Stop 150-21, Pasadena, CA 91125.

The rotational spectrum of methanol is characterized by an asymmetric top overall rotation coupled with an intermediate barrier three fold hindered rotation. The spectrum progresses from a torsionally modified asymmetric top in the v_{t} = 0 state to a rotationally modified torsion by the v_{t} = 2 state. The v_{t} = 3 state lies well above the torsional barrier and in close proximity to the CO stretch vibration. The v_{t} = 3 microwave spectrum has been assigned up to J = 34 in over 1 THz of frequency measured spectra. In spite of the success in assigning the a-type R-branches the b-type spectra is surprisingly weak suggesting that the torsion significantly modifies the intensities of the few K levels in close enough proximity to have transitions below 2.5 THz.